• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 15:42:10 on October 1, 1998

    Entry number 2069

    keyword=Beam Time Accounting
    Hall A

    10/1/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: J. Gomez
    Shift Workers: S. Jensen, M. Liang, R. Ransome

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0.9 0.6 0 0 0 0 1 1 3.5
    Beam Available
    but Not in Use
    0 0.4 1 1 1 0.8 0 0 4.2
    Beam Not Available
    or Unacceptable
    0.1 0 0 0 0 0.2 0 0 0.3
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 0.6 0 0 0 0.2 1 1 3.8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0.4 1 1 1 0.8 0 04.2
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0.4 1 0 1 0.8 0 0 3.2

    Comments: 7-9:36 Data taking at parallel kin 1.484 Gev/c 9:36 - 13:45 Planned Hall access; (i) connect interlock - turn lasers off when heater air flow stops (ii) Check potential cause of Q1H trips due to He Level (iii) install remote threshold adjustment of VDC hadron Points (i) and (ii) got accomplished (see Elog). Scintillator HV (electron arm) were also adjusted - see logbook page 146 14:00-16:00 Moller run

    Data Entry by: J Gomez